OpenStudy (anonymous):

\int_{-\Pi/2 }^{\Pi/2} \frac{cos^2x}{5cos^2x + 4cosx|cosx| + sin2x}

OpenStudy (anonymous):

\[\int\limits_{-\Pi/2 }^{\Pi/2} \frac{\cos^2x}{5\cos^2x + 4cosx|cosx| + \sin2x}\]

OpenStudy (anonymous):

if you are lucky this is even and the integral is 0. if it is not even good luck!

OpenStudy (anonymous):

nope undefined

OpenStudy (anonymous):

So it can't be solved?